If A and B are two independent events together with ?(? ??? ?) = 0.6 and ?(?) = 0.7, calculate ?(?).

If A and B are two independent events and ?(?) = 1 2 and ?(?) = 1 5 , find ?(? ??? ?).

Given that ?(? ?? ?) = 0.86, ?(?) = 0.79, and ?(? ??? ?) = 0.41, find ?(?).
Answer #1

Solution :

If A and B are independent then ,

P(A and B) = P(A) * P(B)

1)

P(B) = P(A and B) / P(A) = 0.6 / 0.7 = 0.8571

2)

P(A and B) = 1/2 * 1/5 = 1 / 10

3)

P(B) = P(A OR B) + ?(? ??? ?) - P(A)

= 0.86 + 0.41 - 0.79

= 0.48